OpenStudy (anonymous):

Evaluate the expression using an appropriate double angle formula. 1/sin(105)cos(105) help!!

jimthompson5910 (jim_thompson5910):

using the formula found on this page http://www.sosmath.com/trig/Trig5/trig5/trig5.html

jimthompson5910 (jim_thompson5910):

look under "double angle formulas"

jimthompson5910 (jim_thompson5910):

we see that sin(2u) = 2sin(u)cos(u) which means that sin(2u) = 2sin(u)cos(u) (1/2)*sin(2u) = sin(u)cos(u) sin(u)cos(u) = (1/2)*sin(2u)

jimthompson5910 (jim_thompson5910):

Let me know if this is enough to get started